1. Find the following: GIVEN: Density of Water: 0.997770 g/cm3 Mass of Water: 24.936 g Mass of Salt (NaCl) Solution: 28.345 g a.) Density of the salt solution: b.) Molarity of the salt in the solution
Answer #1

We know, density = mass/volume.

Hence volume = mass/density = 24.936/0.99777 = 24.992 cm3.

If we consider the salt solution to have the same volume, then we can say:

(a) Density of the salt solution = mass/volume = 28.345/24.992 = 1.134 g/cm3.

(b) Change in density for the salt solution from pure water = 1.134 - 0.99777 = 0.1364 g/cm3.

Hence we can say that per cm3 water will have 0.1364 g more mass. This mass will come from the NaCl added.

Molecular weight of NaCl is 58.44 g/mol. Hence 0.1364 g of NaCl will have = (0.1364/58.44) = 0.00233 moles.

So we have 0.00233 moles of NaCl per cm3 of water. Since 1 cm3 is 0.001 liter, we can write:

Molarity of the salt in solution = 0.00233 / 0.001 = 2.33 M.